OpenStudy (anonymous):

the average speed of 20km/hr,16km/hr,12km/hr?

OpenStudy (anonymous):

20*16*12/20+16+12=80

OpenStudy (anonymous):

the average is 16

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

20 + 16 + 12 = 48/3 = 16

OpenStudy (anonymous):

well i got 3 because there are 3 numbers

OpenStudy (anonymous):

nope

ganeshie8 (ganeshie8):

I think you need to do below : average speed = \[\large \dfrac{3(20*16*12)}{12*16 + 16*12 + 12*20}\]

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

say you have travelled first 1/3rd of distance at 20km/hr next 1/3rd of distance at 16km/hr and the last 1/3rd of distance at 12 km/hr

ganeshie8 (ganeshie8):

then, total distance = d/3 + d/3 + d/3 = d

ganeshie8 (ganeshie8):

time taken while covering first 1/3rd of distance = \(\large \dfrac{d/3}{20}\) time taken while covering second1/3rd of distance = \(\large \dfrac{d/3}{16}\) time taken while covering last /3rd of distance = \(\large \dfrac{d/3}{12}\)

OpenStudy (anonymous):

yes exACTLY @ganeshie8

ganeshie8 (ganeshie8):

so total time taken for covering a distance of "d" = \(\large \dfrac{d/3}{20} + \dfrac{d/3}{16} + \dfrac{d/3}{12} \)

ganeshie8 (ganeshie8):

average speed = (total distance)/(total time) = \(\large \dfrac{d}{\dfrac{d/3}{20} + \dfrac{d/3}{16} + \dfrac{d/3}{12}} \)

ganeshie8 (ganeshie8):

= \(\large \dfrac{1}{\dfrac{1/3}{20} + \dfrac{1/3}{16} + \dfrac{1/3}{12}} \)

ganeshie8 (ganeshie8):

= \(\large \dfrac{3}{\dfrac{1}{20} + \dfrac{1}{16} + \dfrac{1}{12}} \)

ganeshie8 (ganeshie8):

you can simplify

OpenStudy (anonymous):

15.31